OpenStudy (anonymous):

Find f^(302)(x) where f(x) = cos(5x). You do not need to simplify your answer.

OpenStudy (tkhunny):

302/4 = 75 r 2 The original equation and it's first 4 derivatives. \(f^{[0]} = \cos(5x)\) \(f^{[1]} = -5\sin(5x)\) \(f^{[2]} = -5^{2}\cos(5x)\) \(f^{[3]} = 5^{3}\sin(5x)\) \(f^{[4]} = 5^{4}\cos(5x)\) Find a pattern and exploit it.